OpenStudy (anonymous):

how to use the sum or difference identity to prove cos (x- pi/2)= sinx I don't remember how to do it please helps

OpenStudy (anonymous):

\[\cos(A - B) = \cos(A)\cos(B) + \sin(A)\sin(B)\]

OpenStudy (anonymous):

Here, A = x and B = pi/2

OpenStudy (anonymous):

Just use it and tell me where you have arrived by using this..

OpenStudy (anonymous):

will do

OpenStudy (anonymous):

I got the answer. I thought I had to use the co function identities, that's why I was confused how to do it. The identity you used I was told is called the compound angle identity . Thank you